OpenStudy (anonymous):

If the degree measure of an angle is 40° less than 3 times the degree measure of its supplementary angle, what is the degree measure of the angle?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i think i have the answer

OpenStudy (anonymous):

Not even funny. Thanks for nothing.

OpenStudy (anonymous):

you didnt get the answer!!???

OpenStudy (anonymous):

Call x is the finding degree Then supplementary angle is 180 - x We have 3(180-x) - x = 40 --> solve for x, we have 540 - 4x = 40 --> x = 500/4 = 25*5 = 125(degree)

OpenStudy (kc_kennylau):

@meghen_elodie Please read the Code of Conduct. \[\Large \text{"Be Nice: I will stay positive, be friendly, and not mean."}\]

OpenStudy (kc_kennylau):

http://openstudy.com/code-of-conduct

OpenStudy (kc_kennylau):

@meshree I am sorry for your unpleasant experience caused by @meghen_elodie

OpenStudy (anonymous):

im sorry that was the wrong link !!1

OpenStudy (anonymous):

Thanks. I appreciate that. I do not appreciate someone trying to "help" me, and they do something completely, irrelevant.